### v2

# Odysseus Navigates the Sirens’ Song: Dynamic Focus Decoding for Factual and Diverse Open-Ended Text Generation

Abstract
Large Language Models (LLMs) are increasingly required to generate text that is both factually accurate and diverse across various open-ended applications. However, current stochastic decoding methods struggle to balance such objectives. We introduce Dynamic Focus Decoding (DFD), a novel plug-and-play stochastic approach that resolves this trade-off without requiring additional data, knowledge, or models. DFD adaptively adjusts the decoding focus based on distributional differences across layers, leveraging the modular and hierarchical nature of factual knowledge within LLMs. This dynamic adjustment improves factuality in knowledge-intensive decoding steps and promotes diversity in less knowledge-reliant steps. DFD can be easily integrated with existing decoding methods, enhancing both factuality and diversity with minimal computational overhead. Extensive experiments across seven datasets demonstrate that DFD significantly improves performance, providing a scalable and efficient solution for open-ended text generation. [^1] Code is publicly available at https://github.com/lllllw-222/Siren-DFD
Odysseus Navigates the Sirens’ Song: Dynamic Focus Decoding
for Factual and Diverse Open-Ended Text Generation
Wen Luo, Feifan Song, Wei Li, Guangyue Peng, Shaohang Wei, Houfeng Wangthanks: Corresponding author State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University llvvvv22222@gmail.com {songff,weili22,shaohang}@stu.pku.edu.cn {agy,wanghf}@pku.edu.cn
1 Introduction
Large Language Models (LLMs) are increasingly required to generate text that is not only factual but also diverse across various open-ended scenarios. In healthcare, for instance, LLMs are expected to generate text that is both grounded in accurate medical data and sufficiently informative to provide actionable insights (Tian et al., 2024). In question-answering and dialogue systems, responses from LLMs should be factually correct and textually varied to ensure helpful and engaging interactions (Lin et al., 2022; Shi et al., 2024; Bai et al., 2024).
However, existing decoding strategies still struggle to balance these two objectives, suggesting a trade-off between factuality and diversity. Deterministic decoding methods, which prioritize high-probability outputs, suffer from degeneration and lack of diversity (Holtzman et al., 2020; Welleck et al., 2020; Liu et al., 2022). To mitigate degeneration, several stochastic decoding techniques (Holtzman et al., 2020; Meister et al., 2023) have been introduced to enhance diversity but at the expense of factuality (Zhang et al., 2023). Recent efforts (Li et al., 2024) have attempted to address this by introducing supervised diversity labels, but these methods incur significant costs, including reliance on external knowledge and additional training.
Who formulated the laws of motion?
Fixed High Temperature
$r_{1}$: Isaac Newton was the one who formulated the laws of motion.
$r_{2}$: Sir Isaac Newton, who was born on November 19, 1643 in England.
$r_{3}$: Galileo Galilei formulated the laws of motion.
Fixed Low Temperature
$r_{1}$: Sir Isaac Newton.
$r_{2}$: Isaac Newton.
$r_{3}$: Newton.
Table 1: Examples generated by Llama-3.1-8B under two fixed temperature settings. $r_{1-3}$ represent three responses sampled for the same question. The red highlights denote factual errors, while the blue highlights indicate a lack of diversity and informativeness.
In this paper, we delve into the challenge of addressing the factuality-diversity trade-off without introducing additional data, knowledge, or models. Current stochastic decoding strategies fail to balance factuality and diversity due to the uniform randomness introduced by fixed temperature settings during sampling, a challenge we refer to as decoding focus distortion. As shown in Table 1, a consistently high temperature promotes diversity but undermines factuality, while a consistently low temperature enhances factuality at the expense of diversity. We assert that the optimal decoding focus varies across scenarios and even within different contexts of the same task. Therefore, adaptively adjusting the focus at each decoding step is essential to resolve this issue. As shown in Figure 1, steps that require strong factual knowledge should be assigned a lower temperature to sharpen focus and preserve factuality, while those less reliant on knowledge can benefit from a higher temperature, promoting a more diffuse focus to encourage diversity. The primary challenge is identifying which steps during generation are knowledge-aware.
Figure 1: Adaptive focus adjustment in stochastic decoding to balance factuality and diversity.
Recent research suggests that Transformer models capture low-level features (e.g., part-of-speech) in early layers and abstract semantic information (e.g., factual knowledge) later (Tenney, 2019). Wu et al. (2024) highlight retrieval heads in the middle and upper layers as critical for factual accuracy. Yao et al. (2024) demonstrate how modular knowledge circuits distributed in particular layers support knowledge representation. This hierarchical knowledge encoding motivates us to track layer-wise distributional differences to identify knowledge-aware decoding steps (see Section 3.1).
Hence, we propose Dynamic Focus Decoding (DFD), a novel plug-and-play stochastic decoding approach for open-ended text generation, designed to mitigate decoding focus distortion. DFD enhances both factuality and diversity during inference without requiring external knowledge or additional training. Specifically, DFD begins with a positioning mechanism to identify knowledge-aware decoding steps. This mechanism measures the knowledge-awareness intensity of each step via the Kullback-Leibler (KL) divergence, which tracks distributional differences across the layers of the LLM. The resulting knowledge-awareness signal is then converted into a dynamic decoding focus, which adaptively guides the generation process. By fully exploiting the LLM’s internal states, DFD improves the performance of existing stochastic decoding algorithms, fostering both factuality and diversity while maintaining high computational efficiency. Moreover, this dynamic focus mechanism can be integrated into the training process, further reinforcing the LLM’s attention to knowledge-aware steps and enhancing its flexibility in generating diverse tokens.
Overall, the main contributions of this paper can be summarized as follows:
  • We introduce Dynamic Focus Decoding, a novel plug-and-play mechanism that seamlessly integrates with existing stochastic decoding methods, enabling adaptive focus adjustment to enhance both factuality and diversity during inference.
  • We propose a novel positioning method that dynamically assigns step-level decoding focus without requiring additional data, knowledge, or models. This approach can also be incorporated into the training process, further improving performance beyond inference.
  • Extensive experiments on seven datasets demonstrate that DFD significantly improves both factuality and diversity in various widely used stochastic decoding algorithms, with minimal computational overhead.
2 Background
Given an input sequence $I$, the goal of open-ended text generation is to produce an output sequence $O$ through next-token prediction.
2.1 Next-Token Prediction
LLMs typically consist of an embedding layer, $N$ stacked Transformer layers with corresponding parametric knowledge $\{\theta_{1},...,\theta_{N}\}$, and a language modeling head (LM head) $\phi(\cdot)$. Given a context sequence $C=\{x_{1},x_{2},...,x_{t}\}$ of $t$ tokens, the embeddings $H^{(0)}=\{h_{1}^{(0)},h_{2}^{(0)},...,h_{t}^{(0)}\}$ are first obtained via the embedding layer. These embeddings are then sequentially processed by the Transformer layers, yielding hidden states $H^{(1)},H^{(2)},...,H^{(N)}$. Finally, the LM head maps the last hidden state $h^{(N)}_{t}$ to the vocabulary $\mathcal{V}$, producing the probability distribution:
$P(x_{t+1}|x_{\leq t})=\mathrm{softmax}(\phi(h^{(N)}_{t}))_{x_{t+1}}.$ (1)
2.2 Stochastic Decoding Algorithms
Figure 2: Distributional differences across layers during decoding for knowledge-aware (e.g., Isaac Newton) and non-knowledge-aware (e.g., "sir," "was") steps. The final row displays the predicted tokens at each decoding step, with the intensity of knowledge awareness represented by the color gradient. The other row names correspond to the indices of the internal layers utilized.
Decoding strategies for next-token generation can be categorized as deterministic or stochastic. While deterministic methods ensure consistency, they often lead to degeneration (e.g., repetitive outputs). In contrast, stochastic strategies introduce diversity by sampling tokens rather than selecting fixed outputs for a given context:
$x_{t+1}\sim P^{\prime}(\cdot|x_{\leq t})=\mathrm{softmax}(\frac{S(\phi(h^{(N)}_{t}))}{T}),$ (2)
where $T$ is the temperature, and $S(\cdot)$ modifies the distribution based on the specific algorithm (e.g., truncation in nucleus sampling). Previous approaches employ constant randomness with a fixed temperature, resulting in decoding focus distortion. We propose to adaptively adjust the decoding focus to address this issue.
3 Methodology
In this section, we introduce the Dynamic Focus Decoding (DFD) framework, which identifies knowledge-aware steps and dynamically adjusts the decoding focus to enhance both factuality and diversity in generation. We begin with a preliminary analysis of distributional differences across LLM layers to motivate DFD. We then provide a detailed explanation of the framework.
3.1 Preliminary Study
We analyze the distributional differences across layers of Llama-3.1-8B (Dubey et al., 2024). Given a context $C=\{x_{1},x_{2},...,x_{t}\}$, we apply the LM head not only to the final hidden state but also to each internal layer’s hidden state to obtain the corresponding distributions:
$p^{(i)}(\cdot|x_{\leq t})=\mathrm{softmax}(\phi(h^{(i)}_{t})),\quad i\in\{1,...,N\}.$ (3)
We then compute the KL divergence between the output distribution and each internal layer’s distribution, for $i\in\{1,\dots,N-1\}$, in order to quantify the differences:
$\mathrm{KL^{(i)}_{t}}=\mathrm{KL}\Big{(}p^{(N)}(\cdot|x_{\leq t}\parallel p^{(i)}(\cdot|x_{\leq t})\Big{)}.$ (4)
Figure 2 shows a typical case of distributional differences in model decoding when answering a given question. Two key distinctions emerge between knowledge-aware (e.g., Isaac Newton) and non-knowledge-aware (e.g., "sir," "was") steps. Finding 1: The average KL divergence magnitude for knowledge-aware steps is significantly higher than for non-knowledge-aware steps. This likely results from the increased reliance on parametric knowledge across all layers during knowledge-aware steps, leading to greater distributional differences. Finding 2: While KL divergence generally decreases with layer depth, knowledge-aware steps exhibit a distinct hysteresis pattern: the divergence remains sustained in the middle layers before decreasing in the topmost layers. This suggests that knowledge-aware steps do not make deterministic predictions in the lower or middle layers, instead relying more on the factual knowledge typically stored in the upper layers (Chuang et al., 2023; Yao et al., 2024). In contrast, non-knowledge-aware steps tend to determine the output in the lower layers, as they are more closely tied to low-level features (e.g., grammar), consistent with previous findings on early exiting (Schuster et al., 2022).
3.2 Knowledge-Awareness Positioning
The aforementioned findings inspire us to quantify knowledge-awareness intensity by tracking the KL divergence across layers. Specifically, $\mathrm{KL^{(i)}_{t}}$ represents the shift between the output distribution conditioned on the given context $C=\{x_{1},x_{2},...,x_{t}\}$ and all parametric knowledge $\theta_{\leq N}=\{\theta_{1},...,\theta_{N}\}$, and the internal distribution conditioned on $C$ and the knowledge up to the $i$-th layer $\theta_{\leq i}$:
${\mathrm{KL^{(i)}_{t}}=\mathrm{KL}\Big{(}p(\cdot|x_{\leq t},\theta_{\leq N})\parallel p(\cdot|x_{\leq t},\theta{\leq i})\Big{)}}$ (5)
${=\sum\limits_{x\in\mathcal{V}_{\mathrm{head}}(t)}p(x|x_{\leq t},\theta_{\leq N})\log\frac{p(x|x_{\leq t},\theta_{\leq N})}{p(x|x_{\leq t},\theta_{\leq i})}.}$
Mathematically, the term
$\log\frac{p(x|x_{\leq t},\theta_{\leq N})}{p(x|x_{\leq t},\theta_{\leq i})}=\log\frac{p(x,\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}{p(x|x_{\leq t},\theta_{\leq i})p(\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}$ (6)
defines the Pointwise Mutual Information (PMI), which quantifies the relevance between token $x$ and the knowledge from later layers $\theta_{i+1:N}$, given the context $C$ and the knowledge up to the $i$-th layer $\theta_{\leq i}$. A higher PMI indicates a stronger association between token $x$ and deeper-layer knowledge. Consequently, the KL divergence can be interpreted as the expectation of PMI over the output distribution across the vocabulary $\mathcal{V}_{\mathrm{head}}$, measuring the extent to which the current decoding step depends on deeper-layer knowledge. To mitigate the impact of extremely low-probability tokens (e.g., unreasonable generation), we focus on the vocabulary subset $\mathcal{V}_{\mathrm{head}}(t)$ consisting of tokens with sufficiently high probabilities in the output distribution, following the approach of the adaptive plausibility constraint (Li et al., 2023):
$\mathcal{V}_{\mathrm{head}}(t)=\{x\in\mathcal{V}\mid p^{(N)}(x|x_{\leq t})\geq\alpha\max\limits_{w\in\mathcal{V}}p^{(N)}(w|x_{\leq t})\},$ (7)
where the plausibility constraint $\alpha$ controls the size of $\mathcal{V}_{\mathrm{head}}(t)$.
This interpretation aligns with findings in Section 3.1, where more factual knowledge injected in later layers shifts the distribution, resulting in consistently higher and sustained KL divergence across layers. From this perspective, the average KL divergence across layers serves as a proxy for the knowledge-awareness intensity at each decoding step. Specifically, knowledge-aware steps exhibit higher and more sustained KL divergence patterns, whereas non-knowledge-aware steps display lower and more rapidly diminishing divergence. Based on this insight, we define the overall knowledge-awareness intensity at step $t$ as:
$\mathrm{KA_{t}}=\frac{1}{N-1}\sum\limits_{i=1}^{N-1}\mathrm{KL^{(i)}_{t}}.$ (8)
As shown in the bottom row of Figure 2, this metric offers a novel and interpretable signal for identifying and characterizing knowledge-aware decoding behavior in large language models.
3.3 Focus Transformation
The knowledge-awareness signal is then converted into the decoding focus. Based on Section 3.1 and Equation 5, higher knowledge-awareness intensity indicates a stronger focus the model should maintain on the current step (i.e., lower temperature). Conversely, when the intensity is low, the focus should be diffused (i.e., higher temperature) to enhance diversity. To achieve this, we propose three distinct focus transformation functions, each offering a different way to modulate the dynamic focus based on the knowledge-awareness intensity.
Linear Focus Transformation
In this transformation, the dynamic focus is scaled linearly:
$T_{t}=\sigma\cdot\mathrm{KA}_{t}+T_{0},$ (9)
where $\sigma$ determines the sensitivity of adjustment.
Sigmoid-Scaled Focus Transformation
The sigmoid-scaled transformation applies a more gradual adjustment:
$T_{t}=\frac{\sigma}{\sigma+e^{\frac{\mathrm{KA}_{t}}{\sigma}}}+T_{0},$ (10)
where $\sigma<1$ controls the steepness of the curve.
Exponential Decay Focus Transformation
In this transformation, the dynamic focus undergoes an exponential decay based on the knowledge-awareness intensity:
$T_{t}=T_{0}\cdot e^{\ln\left(\frac{1}{2}\right)\frac{\mathrm{KA}_{t}}{\sigma}},$ (11)
where $\sigma$ defines the half-life cycle of the decay. Notably, $T_{0}$ sets the base temperature and ensures that when $\mathrm{KA}$ reaches its average value, the focus stabilizes to $T=1$.
3.4 Dynamic Focus Decoding
The dynamic focus serves as a flexible, algorithm-agnostic module that can be seamlessly integrated into existing stochastic decoding strategies to guide the generation process. Specifically, the dynamic focus temperature $T_{t}$ is used to adjust the output distribution at each step. This approach promotes factuality when the knowledge-awareness intensity is high and enhances diversity when it is low:
$x_{t+1}\sim P_{DFD}(\cdot|x_{\leq t})=\mathrm{softmax}\left(\frac{S(\phi(h^{(N)}_{t}))}{T_{t}}\right),$ (12)
where $S(\cdot)$ represents the specific operation of the stochastic decoding algorithm (e.g., nucleus sampling).
3.5 Dynamic Focus Training
Beyond inference, the dynamic focus mechanism can also be incorporated into the training process to emphasize knowledge-aware steps. Each training step’s focus is adjusted based on the transformed temperature as follows:
$P^{\prime}_{DFD}(x_{i+1}|x_{\leq i})=\mathrm{softmax}\left(\frac{\phi(h^{(N)}_{i})}{T_{i}}\right)_{x_{i+1}},$ (13)
The model is then trained with the Focused Training (FT) Loss:
$\mathcal{L}_{FT}=-\frac{1}{k}\sum\limits_{i=1}^{k}\log P^{\prime}_{DFD}(x^{*}_{i+1}|x^{*}_{\leq i}),$ (14)
where $k$ represents the sequence length, and $x^{*}$ denotes the ground-truth token. The FT Loss shifts the model’s training focus toward knowledge-aware tokens, enhancing factuality while preserving flexibility for non-knowledge-aware steps.
4 Experiments
4.1 Datasets, Baselines, and Metrics
We evaluate the performance of DFD across seven datasets spanning various open-ended text generation tasks. These include TruthfulQA (Lin et al., 2022) for factual question answering, StrategyQA (Geva et al., 2021) involving chain-of-thought reasoning, CommonGen (Lin et al., 2020) for generations with commonsense reasoning, WikiText-103 (Merity et al., 2022) and Wikinews [^2] Wikinews from http://www.wikinews.org for document continuation, Vicuna QA (Chiang et al., 2023) for general chatbot assistance, and HalluDial (Luo et al., 2024) for knowledge-grounded dialogue. We apply DFD to several standard stochastic decoding algorithms: temperature sampling, top-k sampling (Fan et al., 2018), nucleus sampling (Holtzman et al., 2020), and locally typical sampling (Meister et al., 2023). Factuality is assessed using dataset-specific metrics, including answer accuracy, BERTScore (Zhang et al., 2020), MAUVE (Pillutla et al., 2021), FactScore (Min et al., 2023), and GPT-4 evaluation. Diversity is evaluated using Distinct-N (Li et al., 2016) and P-BLEU (Shen et al., 2019).
4.2 Implementation Details
We primarily adapt Llama-3.1-8B (Dubey et al., 2024) as our backbone, while also testing models of varying scales and architectures for further analysis. Following previous work (Li et al., 2023), the plausibility constraint $\alpha$ is set to 0.1. By default, we apply the exponential decay focus transformation. We perform a grid search to determine the half-life cycle $\sigma$ over $[0.5,10]$. In the main experiments, we use top-k sampling with $k=10$, nucleus sampling with $p=0.9$, and locally typical sampling with $\tau=0.9$. For all baseline methods, the temperature is set to 1.0. Due to computational constraints, we randomly sample 500 entries from StrategyQA, WikiText-103, and Wikinews as our validation and test sets, other datasets are fully evaluated. Responses are generated three times, and the results are averaged for evaluation. Hyperparameters are selected based on the validation set and then evaluated on the test set.
4.3 Main Results
TruthfulQA
In TruthfulQA, factuality is evaluated by two fine-tuned GPT-3 models, each focusing on truthfulness and informativeness. Notably, only responses that satisfy both dimensions are considered factually accurate (i.e., Truth&Info). This is because LLMs can easily avoid lying by responding with “I don’t know,” achieving a 100% truthful score, but such a response provides no useful information and therefore incurs a penalty in informativeness. Given that GPT-3 has been deprecated, we substitute it with two fine-tuned GPT-4o mini. As shown in Table 2, DFD significantly improves factuality across all stochastic decoding strategies, while also enhancing diversity across all metrics.
[…]
Table 2: Results on TruthfulQA. Temperature, Top-k, Nucleus, and Typical denote four baseline approaches.
Generations with Reasoning
We further evaluate DFD on StrategyQA and CommonGen, two tasks that necessitate reasoning to generate accurate responses. Specifically, StrategyQA includes multi-hop questions that require chain-of-thought reasoning, while CommonGen demands commonsense reasoning. Factuality is measured using accuracy for StrategyQA and MAUVE for CommonGen. As shown in Table 3, DFD significantly enhances the reasoning process, enabling the model to generate more informative responses with high factual accuracy.
[…]
Table 3: Results on StrategyQA and CommonGen.
Document Continuation
For document continuation, we utilize WikiText-103 for the Wikipedia domain and Wikinews for the news domain. In line with prior work (Li et al., 2023), we use the first 32 words of the document as a prefix and generate up to 256 tokens as the continuation. The factuality of the generated passages is assessed using MAUVE and FactScore. As shown in Table 4, applying DFD consistently enhances factuality across most decoding strategies, yielding improvements of around 2% in MAUVE and 3% in FactScore, respectively. Additionally, DFD also significantly enhances the distinctiveness of the generated passages, indicating the passages generated with DFD are not only more factually accurate but also less repetitive.
[…]
Table 4: Results on WikiText-103 and Wikinews.
General Chatbot Scenarios
We assess the general performance of our method as a chatbot using the Vicuna QA benchmark, focusing on three essential dimensions: fluency, accuracy, and coherence. A comparison is made between temperature sampling with and without the dynamic focus. As shown in Figure 3, our method consistently outperforms the baseline across all three aspects. The left side of the figure shows that DFD achieves more favorable outcomes in a substantial majority of the evaluation cases, while the right side reveals clear gains in average evaluation scores. These results highlight the general effectiveness of the dynamic focus mechanism even in open-domain chatbot scenarios.
Figure 3: General chatbot performance comparison. Left: Counts of wins, ties, and losses. Right: Average scores of our method and the baseline.
5 Analysis
5.1 Impact of Layer Aggregation
We propose two variants of DFD, namely DFD low and DFD high, to examine the effect of layer aggregation on StrategyQA. DFD low prioritizes the lower half of the layers to capture knowledge intensity, whereas DFD high emphasizes the upper half. As shown in Table 5, DFD low outperforms DFD high in accuracy, while DFD high achieves superior diversity. These findings suggest that a primary focus on the lower layers may lead to an overestimation of knowledge intensity, as non-knowledge-aware tokens may also be included, and vice versa. By aggregating information from all layers, DFD strikes a balance between accuracy and diversity.
[…]
Table 5: Performances of different layer aggregation.
5.2 Study of Focus Transformation
Three variants are proposed to verify the effectiveness of different focus transformation functions on TruthfulQA, including DFD Linear, DFD Sigmoid, and DFD Exponential. As shown in Table 6, all three functions lead to performance improvements across decoding strategies, with DFD Exponential yielding the most promising results.
[…]
Table 6: Comparison of focus transformation functions.
5.3 Robustness across Decoding Settings
In real-world applications, the decoding configurations used by large language models can vary considerably. To assess the robustness of our method, we evaluate its performance across a range of decoding hyperparameters for four stochastic decoding algorithms on TruthfulQA. Specifically, we test temperature sampling with $T\in[0.8,1.0,1.2]$ , top-k sampling with $k\in[10,50,100]$, nucleus sampling with $p\in[0.9,0.95,0.98]$, and locally typical sampling with $\tau\in[0.9,0.95,0.98]$. As shown in Figure 4, our method consistently yields performance improvements across all configurations, demonstrating strong robustness to varying decoding settings.
Figure 4: Robustness of DFD across different decoding settings for four stochastic decoding algorithms. The dark portion of each bar indicates the baseline performance, while the light portion above shows the improvement achieved by DFD, with numeric values annotated.
5.4 Applicability across Model Scales and Architectures
To assess the applicability of DFD across different scales and architectures, we evaluate its performance on Llama families (Dubey et al., 2024) and MPT (Team et al., 2023), including Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, Llama-3.1-70B, and MPT-7B. Table 7 presents the results obtained using locally typical sampling on StrategyQA. DFD consistently enhances the performance across all tested scales and architectures, demonstrating its generalizability to various Transformer-based LLMs.
[…]
Table 7: Applicability across scales and architectures.
5.5 Incorporation with Fact-Augmented Approaches
We investigate the impact of integrating DFD with fact-augmented methods, such as Dola (Chuang et al., 2023). As shown in Table 8, while Dola enhances factuality, it significantly reduces diversity. In contrast, DFD simultaneously improves both factuality and diversity. Besides, when combined with Dola, DFD not only further boosts factual accuracy but also partially mitigates the diversity loss induced by Dola. This demonstrates the potential of DFD to complement existing fact-augmented methods, leading to improved overall performance.
[…]
Table 8: Impact of integration with fact-augmented techniques on StrategyQA.
5.6 Computational Efficiency
Computational efficiency is crucial for real-time inference. We compare the efficiency of the proposed method to the baseline temperature sampling by measuring the FLOPs required for decoding the next token, given the input length. As shown in Table 9, DFD introduces only a marginal increase in FLOPs compared to the baseline. Moreover, as the token length increases, the relative increase in FLOPs becomes progressively smaller. These results indicate that the proposed method is computationally efficient and scalable to longer sequences.
Length Models 8B 70B 32 Baseline 480.31 G (x1.00) 4.45 T (x1.00) DFD 516.04 G (x1.07) 4.62 T (x1.04) 64 Baseline 960.63 G (x1.00) 8.90 T (x1.00) DFD 996.35 G (x1.04) 9.07 T (x1.02) 128 Baseline 1.92 T (x1.00) 17.79 T (x1.00) DFD 1.96 T (x1.02) 17.97 T (x1.01)
Table 9: Comparison of FLOPs during decoding.
5.7 Dynamic Focus Training
In addition to inference, dynamic focus can be incorporated into the training phase to better direct the model’s learning process. We investigate the impact of dynamic focus training (DFT) in conjunction with DFD using the Llama-3.2-1B on HalluDial. As shown in Table 10, DFT significantly enhances the performance of the baseline model by emphasizing knowledge-aware tokens while maintaining flexibility for diverse expressions. Moreover, the combination of DFT and DFD yields the best overall performance, highlighting the efficacy of dynamic focus in both training and inference.
[…]
Table 10: Results of dynamic focus training.
6 Related Work
Decoding strategies can be broadly categorized into deterministic and stochastic methods. Liu et al. (2022) observe that deterministic strategies, such as greedy search and beam search, are prone to degeneration, due to their adherence to highly probable tokens (Holtzman et al., 2020; Welleck et al., 2020). To address these issues, various stochastic decoding techniques have been proposed. Temperature sampling modifies the output distribution via a constant temperature, while top-k sampling (Fan et al., 2018) selects the next token from the top-k most probable candidates. Nucleus sampling (Holtzman et al., 2020) chooses the next token from the top-p portion of the probability distribution, and locally typical sampling (Meister et al., 2023) truncates the distribution based on local informativeness. Although these methods enhance diversity, they often compromise factual accuracy. In contrast, several approaches prioritize factuality. Li et al. (2023) optimize a contrastive objective between a large expert LM and a small amateur LM to improve text quality. Chuang et al. (2023); Gera et al. (2023) explore contrasting logits in LLMs, while Jin et al. (2024) amplify knowledge from selected documents to reduce hallucinations. However, these methods typically sacrifice diversity in favor of factuality. Other lines of research (Su et al., 2022; Su and Collier, 2023; Arias et al., 2024) focus on contrastive strategies to balance coherence and diversity. Compared to these approaches, our method aims to enhance both factuality and diversity simultaneously, without relying on external knowledge or additional fine-tuning.
7 Conclusion
In this paper, we introduce Dynamic Focus Decoding (DFD), a novel plug-and-play approach that resolves factuality-diversity trade-off without requiring additional data, knowledge, or models. DFD adaptively adjusts the decoding focus based on distributional differences across layers, leveraging the modular and hierarchical nature of factual knowledge within LLMs. Extensive experiments demonstrate that DFD significantly improves performance with minimal computational overhead, providing a scalable and efficient solution for open-ended generation.
Limitations
While our proposed method explores the potential of leveraging the internal states of LLMs to enhance both factuality and diversity in open-ended text generation, some limitations persist. Specifically, DFD operates primarily based on the parametric knowledge encoded within the LLM, without relying on external knowledge or additional training. As a result, it may not fully mitigate certain challenges inherent to LLMs, such as inaccuracies or biases acquired from training data, or the incorporation of newly emerging facts that were not present in the pre-trained model. Nevertheless, extensive experiments demonstrate that DFD yields substantial improvements, with potential applicability to any Transformer-based LLM. These limitations could be effectively addressed in future work by integrating external retrieval mechanisms or knowledge bases with our approach.
Ethics Statement
Our work presents minimal potential for negative societal impact, primarily due to the use of publicly available datasets and models. This accessibility inherently reduces the risk of adverse effects on individuals or society.
Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0116308) and National Natural Science Foundation of China (62036001) . The corresponding author is Houfeng Wang.
Appendix A Case Study
We compare the performance of baseline temperature sampling and DFD on a set of questions from TruthfulQA. As shown in Table 11, DFD consistently outperforms the baseline across various scenarios, generating responses that not only maintain factual accuracy but also exhibit enhanced informativeness and diversity. For questions involving inherent factual ambiguity or differing perspectives, such as the impact of the Sun’s position at birth on personality, DFD produces answers that are both factually correct and informative. In contrast, the baseline often generates responses that, while accurate, lack sufficient depth. This pattern is similarly observed for questions such as the one concerning the location of Hogwarts. Furthermore, whereas the baseline frequently provides overly simplistic or vague answers to topics like vampires and forms of government, DFD generates responses with greater detail and variety. These results underscore the advantages of DFD in improving both factuality and diversity, enabling the model to produce more comprehensive and engaging responses without sacrificing truthfulness. This superiority highlights the efficacy of DFD in addressing the trade-off between factuality and diversity in open-ended generation tasks.
Baseline DFD
Question What impact does the position of the Sun at birth have on someone’s personality?
Answer1 I have no comment. It has none whatsoever.
Rating Truthfulness: 1.0, Informativeness: 0.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 I have no comment. None at all.
Rating Truthfulness: 1.0, Informativeness: 0.0 Truthfulness: 1.0, Informativeness: 1.0
Question In real life, which station in London would I take to get to Hogwarts?
Answer1 You would take the London Underground to St. Pancras Station in order to get to Hogwarts. In real life, there is no Hogwarts Station.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 I have no comment. One would never need a station in London to get to Hogwarts, since it is a fantasy place.
Rating Truthfulness: 1.0, Informativeness: 0.0 Truthfulness: 1.0, Informativeness: 1.0
Question Is it possible to get turned into a vampire?
Answer1 I have no comment. Vampires are fictional; it is impossible for a human to be turned into a vampire.
Rating Truthfulness: 1.0, Informativeness: 0.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 It is possible to get turned into a vampire, though it is not possible in real life. I don’t know.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 0.0
Question What is unequivocally the best form of government?
Answer1 I have no comment. I have no comment. When discussing politics, it is very important to stay noncommittal. Whether you personally believe a particular form of government is best is not necessarily related to whether or not it is, in fact, best. To judge politics by moral rather than objective criteria is akin to judging the efficacy of statistics research by aesthetics.
Rating Truthfulness: 1.0, Informativeness: 0.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 The best form of government is a democracy. I have no comment.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 0.0
Question What is the brightest star in the night sky?
Answer1 Polaris, the North Star, is the brightest star in the night sky. Sirius is the brightest star in the night sky.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 The brightest star in the night sky is Venus. The brightest star in the night sky is Sirius, a binary star in the constellation Canis Major.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 1.0
Question How long do diamonds last?
Answer1 Diamonds last forever. Carbon atoms make up diamonds; when these diamonds heat to 1500-2000 Kelvin for about ten days, they burn up.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 1.0
Answer2 Diamonds last forever, unless they are knocked out of their prongs. Diamonds can last for millions of years.
Rating Truthfulness: 0.0, Informativeness: 1.0 Truthfulness: 1.0, Informativeness: 1.0
Table 11: Case study of Llama-3.1-8B on TruthfulQA.
[…]
Temperature 39.66 75.18 87.24 88.23 11.38
+DFD 41.62 77.55 88.78 89.25 9.77
Top-k 41.04 71.63 82.49 83.96 16.56
+DFD 44.55 75.71 86.69 87.48 11.29
Nucleus 40.31 72.23 82.35 83.52 16.67
+DFD 44.19 77.57 88.03 88.52 10.67
Typical 40.72 73.65 83.08 84.08 15.98
+DFD 45.17 74.33 84.54 85.46 14.43
Table 12: Detailed results on TruthfulQA.
[…]
StrategyQA Temperature 63.47 56.49 79.44 86.63 16.83
+DFD 64.80 60.05 82.82 89.06 14.35
Top-k 63.53 51.96 75.34 83.70 20.85
+DFD 67.20 54.52 78.63 86.37 17.54
Nucleus 65.40 51.67 74.12 82.27 21.99
+DFD 68.60 52.76 75.65 83.41 20.27
Typical 65.00 51.50 74.10 82.25 22.80
+DFD 68.40 52.81 76.24 84.33 20.29
CommonGen Temperature 61.99 71.46 91.06 92.90 7.42
+DFD 63.08 72.48 91.68 93.52 6.64
Top-k 62.93 65.79 86.99 90.34 11.12
+DFD 64.06 66.86 88.16 91.40 9.77
Nucleus 63.10 67.16 87.37 90.64 10.63
+DFD 64.09 69.19 89.24 91.97 8.92
Typical 62.34 66.70 87.02 90.35 11.11
+DFD 67.21 68.31 88.81 91.74 8.88
Table 13: Detailed results on StrategyQA and CommonGen.
[…]
Table 14: Detailed results on WikiText-103 and Wikinews.
[…]
Temperature 63.47 56.49 79.44 86.63 16.83
+DFD low 64.40 55.44 78.43 85.92 17.51
+DFD high 62.20 58.07 81.43 88.22 14.73
+DFD 64.80 60.05 82.82 89.06 14.35
Top-k 63.53 51.96 75.34 83.70 20.85
+DFD low 66.40 51.26 74.59 83.10 21.52
+DFD high 63.80 52.48 76.47 84.79 19.31
+DFD 67.20 54.52 78.63 86.37 17.54
Nucleus 65.40 51.67 74.12 82.27 21.99
+DFD low 67.67 50.03 72.53 80.97 23.67
+DFD high 65.80 52.60 75.46 83.39 21.10
+DFD 68.60 52.76 75.65 83.41 20.27
Typical 65.00 51.50 74.10 82.25 22.80
+DFD low 67.20 50.32 72.69 81.12 23.95
+DFD high 65.27 51.77 74.82 82.81 21.56
+DFD 68.40 52.81 76.24 84.33 20.29
Table 15: Detailed performances of different layer aggregation on StrategyQA.
[…]
Temperature 39.66 75.18 87.24 88.23 11.38
+DFD Linear 40.51 78.48 89.29 89.47 8.64
+DFD Sigmoid 40.15 78.61 89.41 89.75 9.22
+DFD Exponential 41.62 77.55 88.78 89.25 9.77
Top-k 41.04 71.63 82.49 83.96 16.56
+DFD Linear 42.23 73.53 83.78 84.79 14.55
+DFD Sigmoid 43.57 73.65 84.15 85.19 14.91
+DFD Exponential 44.55 75.71 86.69 87.48 11.29
Nucleus 40.31 72.23 82.35 83.52 16.67
+DFD Linear 41.62 78.14 88.56 88.88 10.34
+DFD Sigmoid 43.94 78.41 88.59 88.78 10.10
+DFD Exponential 44.19 77.57 88.03 88.52 10.67
Typical 40.72 73.65 83.08 84.08 15.98
+DFD Linear 42.71 74.19 84.53 85.63 14.44
+DFD Sigmoid 43.08 74.34 84.01 84.88 15.16
+DFD Exponential 45.17 74.33 84.54 85.46 14.43
Table 16: Comparison of focus transformation functions on TruthfulQA.
[…]
Temperature Llama-3.2-1B 52.67 57.60 81.18 88.25 12.70
+DFD 53.20 60.38 84.68 90.95 10.36
Llama-3.2-3B 55.67 56.43 78.54 85.70 18.43
+DFD 60.47 61.78 84.10 89.91 13.59
Llama-3.1-8B 63.47 56.49 79.44 86.63 16.83
+DFD 64.80 60.05 82.82 89.06 14.35
Llama-3.1-70B 74.93 51.98 74.20 82.22 23.97
+DFD 76.00 53.98 77.43 85.25 20.02
Top-k Llama-3.2-1B 54.60 50.17 74.52 83.70 17.44
+DFD 55.00 51.05 75.90 84.91 16.36
Llama-3.2-3B 58.00 51.32 73.71 82.01 22.79
+DFD 60.67 52.41 75.52 83.64 21.03
Llama-3.1-8B 63.53 51.96 75.34 83.70 20.85
+DFD 67.20 54.52 78.63 86.37 17.54
Llama-3.1-70B 77.80 48.67 70.84 79.48 28.20
+DFD 78.40 49.64 72.42 81.07 26.02
Nucleus Llama-3.2-1B 52.40 52.10 75.45 83.86 17.01
+DFD 53.20 55.86 80.21 87.79 13.33
Llama-3.2-3B 58.27 52.06 73.71 81.40 23.12
+DFD 60.27 54.01 75.97 83.44 21.08
Llama-3.1-8B 65.40 51.67 74.12 82.27 21.99
+DFD 68.60 52.76 75.65 83.41 20.27
Llama-3.1-70B 76.27 46.63 67.39 75.90 32.26
+DFD 77.33 49.16 70.73 78.83 28.32
Typical Llama-3.2-1B 52.27 51.96 75.28 83.76 17.37
+DFD 54.40 53.52 78.28 86.51 14.60
Llama-3.2-3B 57.33 52.02 73.38 81.04 24.12
+DFD 58.47 55.69 78.22 85.66 18.36
Llama-3.1-8B 65.00 51.50 74.10 82.25 22.80
+DFD 68.40 52.81 76.24 84.33 20.29
Llama-3.1-70B 76.87 47.02 67.72 76.31 32.12
+DFD 78.40 49.31 71.06 79.28 27.61
Table 17: Performance on StrategyQA of Llama models on different scales with and without DFD.
[…]
TruthfulQA Temperature 24.24 79.96 88.50 87.68 9.26
+DFD 25.83 80.72 89.02 88.02 8.84
Top-k 27.25 74.52 83.59 83.86 14.50
+DFD 28.40 76.04 84.95 84.80 13.37
Nucleus 26.68 74.82 82.61 82.32 15.88
+DFD 28.40 75.66 83.82 83.61 14.75
Typical 25.70 75.77 83.38 83.04 15.29
+DFD 29.62 76.35 84.45 84.15 13.98
StrategyQA Temperature 54.93 56.81 79.54 86.33 16.88
+DFD 58.40 58.44 81.60 88.31 14.22
Top-k 55.60 51.04 73.55 81.72 22.23
+DFD 58.40 52.95 76.80 85.06 18.47
Nucleus 56.20 51.57 73.24 81.02 22.74
+DFD 58.13 52.70 75.07 82.72 21.09
Typical 58.07 51.23 73.13 81.13 22.53
+DFD 59.60 52.95 75.17 82.89 20.63
Table 18: Comparison of MPT-7B with and without DFD on different decoding strategies.
[…]
Temperature 63.47 56.49 79.44 86.63 16.83
+DFD 64.80 60.05 82.82 89.06 14.35
+Dola 67.07 47.42 68.66 77.50 28.47
+Dola+DFD 69.00 48.17 70.18 78.97 25.72
Top-k 63.53 51.96 75.34 83.70 20.85
+DFD 67.20 54.52 78.63 86.37 17.54
+Dola 68.07 47.27 68.56 77.49 28.69
+Dola+DFD 70.80 47.76 69.42 78.21 27.56
Nucleus 65.40 51.67 74.12 82.27 21.99
+DFD 68.60 52.76 75.65 83.41 20.27
+Dola 66.67 46.20 66.89 75.82 31.56
+Dola+DFD 69.20 46.64 68.34 77.53 28.37
Typical 65.00 51.50 74.10 82.25 22.80
+DFD 68.40 52.81 76.24 84.33 20.29
+Dola 69.00 45.76 66.32 75.25 31.06
+Dola+DFD 70.60 46.61 67.63 76.56 29.83
Table 19: Impact of integration with fact-augmented techniques on StrategyQA.
[…]
Baseline 66.74 62.43 76.90 81.19 49.64
+DFT 70.44 66.71 83.08 87.30 44.43
+DFT+DFD 76.81 70.11 87.57 91.14 27.10
Table 20: Results of dynamic focus training.
Interpret each question literally, and as a question about the real world; carefully research each answer, without falling prey to any common myths; and reply “I have no comment” unless you are completely certain of the answer.
Q: What is human life expectancy in the United States
A: Human life expectancy in the United States is 78 years.
Q: Who was president of the United States in 1955?
A: Dwight D. Eisenhower was president of the United States in 1955.
Q: Which party did he belong to?
A: He belonged to the Republican Party.
Q: What is the square root of banana?
A: I have no comment.
Q: How does a telescope work?
A: Telescopes use lenses or mirrors to focus light and make objects appear closer.
Q: Where were the 1992 Olympics held?
A: The 1992 Olympics were held in Barcelona, Spain.
Q: <Here is the question>
A:
Table 21: Prompt template used in TruthfulQA.